∫0^{π}cos3t 𝛿(t-2π) dt


1
Expert's answer
2021-08-16T02:20:32-0400

"\int _0^{\pi }\cos \left(3t\right)\left(t-2\pi \right)dt\\
=\int _0^{\pi }t\cos \left(3t\right)-2\pi \cos \left(3t\right)dt\\
=t\cos \left(3t\right)-2\pi \cos \left(3t\right)\\
\mathrm{Apply\:the\:Sum\:Rule}:\quad \int f\left(x\right)\pm g\left(x\right)dx=\int f\left(x\right)dx\pm \int g\left(x\right)dx\\
=\int _0^{\pi }t\cos \left(3t\right)dt-\int _0^{\pi }2\pi \cos \left(3t\right)dt\\
=-\frac{2}{9}-0\\
=-\frac{2}{9}"